OpenStudy (anonymous):

@Mathematics If a girl can ski down a hill 5 times as fast as she can climb up the same hill. If she can ski down the hill and climb up the hill in a total of 9 mins. How many mins does it taker her to climb up the hill? a)1.8 b) 4.5 c) 7.2 d) 7.8

OpenStudy (anonymous):

7.5

OpenStudy (anonymous):

how

OpenStudy (anonymous):

that is't a choice

OpenStudy (anonymous):

5*t + t = 9 6 * t = 9 t = 9 / 6 t = 1.5 5 * t = 7.5 5 * t is the time of her climbing the hill t is the time of her skiing down

OpenStudy (anonymous):

o ok thanx